# GEMINUS: Dual-aware Global and Scene-Adaptive Mixture-of-Experts for End-to-End Autonomous Driving

### v3

@@ Abstract
End-to-end autonomous driving requires adaptive and robust handling of complex and diverse traffic environments. However, prevalent single-mode planning methods attempt to learn an overall policy while struggling to acquire diversified driving skills to handle diverse scenarios. Therefore, this paper proposes GEMINUS, a Mixture-of-Experts end-to-end autonomous driving framework featuring a Global Expert, a Scene-Adaptive Experts Group, and equipped with a Dual-aware Router. Specifically, the Global Expert is trained on the overall dataset, possessing robust performance. The Scene-Adaptive Experts are trained on corresponding scene subsets, achieving adaptive performance. The Dual-aware Router simultaneously considers scenario-level features and routing uncertainty to dynamically activate expert modules. Through the effective coupling of the Global Expert and the Scene-Adaptive Experts Group via the Dual-aware Router, GEMINUS achieves adaptive and robust performance in diverse scenarios. GEMINUS outperforms existing methods in the Bench2Drive closed-loop benchmark and achieves state-of-the-art performance in Driving Score and Success Rate, even with only monocular vision input. Furthermore, ablation studies demonstrate significant improvements over the original single-expert baseline: 7.67% in Driving Score, 22.06% in Success Rate, and 19.41% in MultiAbility-Mean. The code will be available at https://github.com/newbrains1/GEMINUS.
Figure 1: The overall architecture of GEMINUS. GEMINUS integrates a Global Expert and a Scene-Adaptive Experts Group. During the training stage, the Global Expert is trained on the overall dataset. Concurrently, each scene-adaptive expert is trained on its respective scene subset guided by scenario-aware routing. In the inference stage, as shown above, features extracted from upstream encoders are processed by the Dual-aware Routing. When scenario uncertainty is below a threshold, the scenario-aware routing activates the highest-scoring adaptive expert; conversely, for high uncertainty, the uncertainty-aware routing activates the Global Expert. This design ensures both adaptive and robust performance across diverse scenarios.
@@ I INTRODUCTION
In recent years, a prominent research direction in autonomous driving has been the development of planning-oriented end-to-end models[1]. In contrast to modular autonomous driving consisting of modular pipelines such as perception, prediction and planning [2, 3, 4], end-to-end methods directly map raw sensor inputs to planned trajectories [5, 6, 7, 8], control signals [9, 10], or a fused output derived from trajectory and control branches [11, 12]. These approaches provide a holistic model for driving, enabling unified optimization towards a global objective, significantly reducing manual engineering efforts, and allowing for the direct use of rich sensor information.
@@
Despite their notable benefits, a persistent limitation of current end-to-end autonomous driving models stems from their global imitation learning on overall training datasets. This approach, typically employing single-mode planning with L2 loss, inherently models the complex output space as a single Gaussian distribution, leading to a tendency towards mode averaging[8, 13]. Consequently, their performance is compromised, as the generated output represents an averaged behavior across diverse scenarios rather than the optimal policy for the current specific scenario. This ultimately restricts the acquisition of diversified driving skills to handle diverse scenarios. Prior approaches employed command-based conditional imitation learning to mitigate mode averaging[9, 14]. However, this approach faced an inherent limitation: solely relying on driving commands is insufficient to distinguish complex scenarios (e.g., an overtaking scenario simultaneously includes turn left, go straight, and turn right commands). Such rigid classification fails to comprehensively consider rich scene information, thus hindering the capture of the diversity of driving skills.
@@
Inspired by the success of Mixture-of-Experts (MoE) architectures in large language models (LLMs) to handle complex data distribution[15], MoE architectures present significant potential for addressing challenges in autonomous driving. By offering a fine-grained scenario adaptation and specialized behavior generation, MoE could mitigate the mode averaging problem and enhance model adaptability in diverse driving scenarios. However, directly transferring generic MoE architectures, designed primarily for static textual data, to autonomous driving reveals an inherent unsuitability. Specifically, they struggle with effective expert specialization due to lack of explicit scenario division, and fail to adequately consider the robustness requirements of autonomous driving.
Therefore, this paper proposes GEMINUS: dual-aware Global and scEne-adaptive MIxture of experts for end-to-end autoNomoUS driving. Specifically, the Global Expert is trained on the overall dataset, possessing robust performance. The Scene-Adaptive Experts are trained on corresponding scene subsets, achieving adaptive performance. The Dual-aware Router simultaneously considers scenario-level features and routing uncertainty to dynamically activate expert modules. Through the effective coupling of the Global Expert and the Scene-Adaptive Experts Group via the Dual-aware Router, GEMINUS simultaneously achieves adaptive and robust performance in diverse scenarios. Our contributions can be summarized as follows:
@@
  • GEMINUS is proposed as a novel Mixture-of-Experts end-to-end autonomous driving framework. This framework effectively integrates a Global Expert and a Scene-Adaptive Experts Group via Dual-aware Router, to simultaneously achieve adaptive performance in feature-distinct scenarios and robust performance in feature-ambiguous scenarios.
  • This paper introduces a Dual-aware Router for end-to-end autonomous driving, uniquely designed with scenario-awareness to identify differences between diverse scenarios, and uncertainty-awareness to model routing uncertainty.
@@
  • To further investigate GEMINUS’s intrinsic routing mechanism, we analyze the impact of the uncertainty threshold on driving performance and examined the router accuracy and expert utilization on the validation set.
@@ II-A End-to-End Autonomous Driving
End-to-end autonomous driving is an emerging paradigm. It directly maps raw sensor inputs to vehicle control actions or planned trajectories. This approach offers significant advantages by simplifying system architecture and mitigating error propagation inherent in cascaded modular pipelines. End-to-end driving policies are predominantly learned using Imitation Learning (IL). This typically involves behavior cloning (BC) to mimic expert demonstrations and capture human-like driving behaviors [5, 6, 7, 8, 9, 10, 11, 12]. Reinforcement Learning (RL) also plays a role. It enables dynamic policy optimization through environmental interaction and reward design [16, 17, 18].
In 2019, CILRS[14] was proposed to obtain control signals by introducing conditional imitation learning. Trajectory-guided Control Prediction (TCP) [11] is a simple and robust monocular vision baseline. This method innovatively integrates trajectory planning and direct control into a unified pipeline for joint learning and predictive fusion. Furthermore, techniques such as TransFuser [5] adopt the Transformer to fuse information from heterogeneous sensors (cameras and LiDAR). Additionally, DriveAdapter [10] employs a student model to learn rich environmental representations from multi-camera information, aiming to overcome the traditional coupling barriers between perception and planning. Beyond fusion, recent innovations also include vectorized scene representations, exemplified by VAD [19], which models driving scenes as fully vectorized elements to improve planning efficiency and robustness. Hydra-MDP [20] explores multimodal planning by distilling knowledge from human and rule-based teachers to generate diverse trajectory candidates. Diffusion models have also emerged as a powerful tool, with DiffusionDrive [21] utilizing truncated diffusion policies to model multi-modal action distributions while simultaneously achieving real-time control. Despite these advancements, existing end-to-end models remain constrained by mode averaging, making it challenging to effectively handle diverse scenarios.
@@ II-B Mixture-of-Experts in Autonomous Driving
MoE architecture has emerged as a significant method for scaling large language models and enhancing task specialization. In Large Language Models, sparse MoE designs boost model capacity and processing efficiency through conditional computation [15]. The strength of MoE lies in its ability to leverage individual experts’ strengths across varying data subsets or tasks, thereby improving overall model performance. In [22], a task-level MoE was applied to multilingual translation, intelligently routing inputs based on linguistic or task identifiers to achieve performance gains and improved inference throughput.
Despite promising results in LLMs, MoE’s application in end-to-end autonomous driving remains underexplored. Some existing studies have explored MoE architectures in autonomous driving for tasks like rare scenario perception [23], long-tailed trajectory prediction[24], domain adaption in different weathers [25], safe trajectory prediction and planning [26], and facilitating generalization of planner [27]. However, these existing approaches have not focused on leveraging MoE to enhance adaptive and robust performance in diverse scenarios.
@@ III METHODOLOGY
Fig. 1 illustrates the overall architecture of GEMINUS end-to-end autonomous driving framework. Informed by certain design philosophies of TCP [11], a single-expert baseline is established. Building upon this baseline, we integrate it with Dual-aware MoE, culminating in GEMINUS end-to-end autonomous driving framework.
@@ III-A Preliminaries
End-to-End Autonomous Driving. The objective of end-to-end autonomous driving is to directly map raw sensor inputs to corresponding trajectories or control actions. In this paper, the raw sensor input $x$ encompasses: a front-facing camera image $i$, the ego-vehicle speed $v$, a high-level navigation command $c$, and a goal point $(x_{g},y_{g})$. The raw sensor inputs are processed by the end-to-end model. Encoders first process these inputs to generate intermediate features. These features are then fed into a trajectory planner. The trajectory planner generates a planned trajectory, comprising waypoints over $K$ steps. This planned trajectory is then fed into a Proportional-Integral-Derivative (PID) controller. The controller subsequently produces the final longitudinal control signals: throttle $\in[0,1]$, brake $\in[0,1]$, and the lateral control signal: steer $\in[0,1]$.
Mixture-of-Experts. MoE architectures offer a principled approach to address the complexities of multimodal data distributions by employing a “divide and conquer” strategy [15]. Introducing the MoE Framework to end-to-end trajectory planning, the overall policy distribution $p_{\theta}(Y\mid X)$ is typically represented as a probabilistic mixture of policy distributions of components $K$, each parameterized by an expert $m_{\theta}(Y\mid Z=k,X)$ and weighted by a gating network $q_{\theta}(Z=k\mid X)$, formalized as:
@@
By flexibly coupling multiple policy distributions, this probabilistic formulation offers a promising framework to effectively model multimodal distributions in end-to-end trajectory planning and tackle the mode averaging problem. Despite potential non-convex optimization challenges in learning such mixture models, deep learning implementations of MoE often simplify this by identifying and assigning the most suitable expert. Notably, models employing a Hard Assignment approach (i.e., selecting a single ”best” expert for a given sample) are highly effective and computationally efficient for multimodal distributions[28]. This is because Hard Assignment directly selects the most matching behavior mode, avoiding the averaging of all experts’ outputs, which further mitigates the mode averaging problem. Building upon these theoretical underpinnings, this paper proposes GEMINUS, a distinctive MoE framework that is specifically tailored for diverse and complex autonomous driving scenarios. At its core, a Dual-aware Router possesses scenario and uncertainty awareness to dynamically activate experts from a Global Expert and a Scene-Adaptive Experts Group. During inference, the Dual-aware Router processes intermediate features $x$ extracted by the encoders. It determines the final output $y$ based on the uncertainty measure $U(x)$ and the scores of scene experts $S_{E}(x)$, formalized as:
@@
Here $x$ is the input feature. $\tau$ denotes a predefined uncertainty threshold. $S$ is the set of all the scene-adaptive experts. When $U(x)$ is low ($U(x)<\tau$), the expert with the highest routing score $S_{E_{i}}(x)$ is selected. This achieves precise and scenario-specific planning in feature-distinct scenarios. In contrast, when the uncertainty of the scenario $U(x)$ is high ($U(x)\geq\tau$), the model selects the Global Expert $f_{\text{global}}(x)$. This ensures robust performance in feature-ambiguous scenarios. Such design allows GEMINUS to effectively avoid the mode averaging problem, thereby achieving adaptive and robust performance in diverse scenarios.
@@ III-B Single-Expert Baseline
Feature Encoders. The image encoder is built on a ResNet34 architecture, pre-trained on ImageNet [29]. This encoder processes the front-facing camera input image and outputs a 1000-dimensional feature embedding vector $I_{\text{feature}}$. Concurrently, a measurement encoder receives a concatenated input $m$ and generates a 128-dimensional measurement feature vector $M_{\text{feature}}$. The input $m$ comprises ego-vehicle speed $v$, a high-level navigation command $c$, and the navigation goal point $(x_{g},y_{g})$. Finally, $I_{\text{feature}}$ and $M_{\text{feature}}$ are concatenated to form combined feature $F$ for the subsequent trajectory planner and router.
Trajectory Planner. The trajectory planner receives the upstream combined feature vector $F$ as input. This input is then passed through a series of linear layers for downsampling, forming a 256-dimensional feature vector $f$. This feature vector $f$ is then fed into a waypoint generator GRU[30]. The GRU model auto-regressively generates future waypoints one by one. This forms a sequence of waypoints $(w_{0},w_{1},...,w_{k})$ for the next $K$= 4 steps. The longitudinal and lateral controllers process these waypoints to generate the final longitudinal control signals (throttle $\in[0,1]$, brake $\in[0,1]$) and the lateral control signal (steer $\in[0,1]$).
Figure 2: Dual-aware MoE vs Vanilla MoE. A vanilla MoE typically tries to balance the load across its experts. In contrast, Dual-aware MoE leverages scenario subset IDs during training. This explicitly directs inputs from specific scenario subsets to their corresponding experts through a scenario-aware routing. During inference, the learned router dynamically activates the most appropriate expert based solely on the input scene features, thus obviating the need for explicit scenario ID. Furthermore, an uncertainty-aware routing is introduced to ensure the model’s robustness and stability. Specifically, when scenario features are ambiguous, preventing the router from effectively distinguishing the true underlying expert, the input is strategically routed to a global expert, ensuring robust and stable model performance.
@@ III-C Scenario-aware Routing Mechanism
Vanilla MoE aims to balance expert usage across GPUs to utilize maximum benefit from features of the inputs. However, this leads to inefficient knowledge sharing among experts when dealing with heterogeneous input distributions. For example, driving policies for a Merging scenario differ significantly from an Emergency Brake scenario. To address this inefficiency and foster specialized knowledge, a scenario-aware routing mechanism is introduced. This mechanism draws inspiration from dataset-aware routing in [31].
@@
$\mathcal{L}_{\text{scenario}}=-\sum_{i=1}^{|E|}\mathbb{1}(h(s_{m})=i)\cdot\log p_{i}(x)$ (3)
$1(\cdot)$ is the indicator function. This loss ensures that all inputs originating from a specific scenario category are primarily dispatched to their designated scene-adaptive expert. By selectively routing inputs based on their scenario feature, this mechanism promotes efficient knowledge specialization within each expert. This enables the model to learn adaptive driving policies.

### v4

@@ Abstract
End-to-end autonomous driving requires adaptive and robust handling of complex and diverse traffic environments. However, prevalent single-mode planning methods attempt to learn an overall policy while struggling to acquire diversified driving skills to handle diverse scenarios. Therefore, this paper proposes GEMINUS, a Mixture-of-Experts end-to-end autonomous driving framework featuring a Global Expert and a Scene-Adaptive Experts Group, equipped with a Dual-aware Router. Specifically, the Global Expert is trained on the overall dataset, possessing robust performance. The Scene-Adaptive Experts are trained on corresponding scene subsets, achieving adaptive performance. The Dual-aware Router simultaneously considers scenario-level features and routing uncertainty to dynamically activate expert modules. Through the effective coupling of the Global Expert and the Scene-Adaptive Experts Group via the Dual-aware Router, GEMINUS achieves both adaptability and robustness across diverse scenarios. GEMINUS outperforms existing methods in the Bench2Drive closed-loop benchmark and achieves state-of-the-art performance in Driving Score and Success Rate, even with only monocular vision input. The code is available at https://github.com/newbrains1/GEMINUS.
Figure 1: The overall architecture of GEMINUS. GEMINUS integrates a Global Expert and a Scene-Adaptive Experts Group. During the training stage, the Global Expert is trained on the overall dataset and each scene-adaptive expert is trained on respective scene subset guided by scenario-aware routing. In the inference stage, as shown above, features extracted from upstream encoders are processed by the Dual-aware Router. When scenario uncertainty is below a threshold, the scenario-aware routing activates the highest-scoring adaptive expert; conversely, for high uncertainty, the uncertainty-aware routing activates the Global Expert. This design ensures both adaptive and robust performance across diverse scenarios.
Figure 2: Dual-aware MoE vs Vanilla MoE. Unlike a vanilla MoE which typically balances expert load, the Dual-aware MoE employs two distinct routing strategies: scenario-aware routing and uncertainty-aware routing. The scenario-aware routing leverages subset IDs during training to guide specific inputs to corresponding experts, enabling the learned router to dynamically activate the most appropriate expert from scene features alone at inference. Furthermore, to ensure robustness in scenarios with ambiguous features, we further introduce an uncertainty-aware routing and a global expert.
@@ I INTRODUCTION
In recent years, a prominent research direction in autonomous driving is the development of planning-oriented end-to-end models[1]. In contrast to modular autonomous driving, which consists of a pipeline of perception, prediction and planning [2, 3, 4], end-to-end methods directly map raw sensor inputs to planned trajectories [5, 6, 7, 8], control signals [9, 10], or a fused output derived from trajectory and control branches [11, 12]. These approaches provide a holistic model for driving, enabling unified optimization towards a global objective, significantly reducing manual engineering efforts, and allowing for the direct use of rich sensor information.
@@
Despite their notable benefits, a persistent limitation of current end-to-end autonomous driving models stems from their global imitation learning on overall training datasets. This approach, typically employing single-mode planning with L2 loss, inherently models the complex output space as a single Gaussian distribution, leading to a tendency toward mode averaging[8, 13]. Consequently, their performance is compromised, as the generated output represents an averaged behavior across diverse scenarios rather than the optimal policy for the current specific scenario. This ultimately restricts the acquisition of diversified driving skills to handle diverse scenarios. Prior approaches employed command-based conditional imitation learning to mitigate mode averaging[9, 14]. However, this approach faced an inherent limitation: solely relying on driving commands is insufficient to distinguish complex scenarios (e.g., an overtaking scenario simultaneously includes turn left, go straight, and turn right commands). Such rigid classification fails to comprehensively consider rich scene information, thus hindering the capture of the diversity of driving skills.
@@
Inspired by the success of Mixture-of-Experts (MoE) architectures in large language models (LLMs) to handle complex data distribution[15], MoE architectures present significant potential for addressing challenges in autonomous driving. By offering a fine-grained scenario adaptation and specialized behavior generation, MoE could mitigate the mode averaging problem and enhance model adaptability in diverse driving scenarios. However, directly transferring generic MoE architectures, designed primarily for static textual data, to autonomous driving reveals an inherent unsuitability. Specifically, they struggle with effective expert specialization due to lack of explicit scenario division and fail to adequately consider the robustness requirements of autonomous driving.
Therefore, this paper proposes GEMINUS: dual-aware Global and scEne-adaptive MIxture of experts for end-to-end autoNomoUS driving. Our contributions can be summarized as follows:
@@
  • We present GEMINUS, a novel MoE framework for end-to-end autonomous driving. This framework adopts a cooperative architecture that partitions the driving policy between a robust Global Expert and a Scene-Adaptive Experts Group to simultaneously achieve adaptive performance in feature-distinct scenarios and robust performance in feature-ambiguous scenarios.
  • We devise a dual-aware Router as the core mechanism of GEMINUS. It is engineered with two complementary awareness capabilities: scenario awareness for precise expert dispatching in distinct situations and uncertainty awareness for ensuring robust fallback in ambiguous conditions.
@@
  • We conducted in-depth analyses of GEMINUS’s intrinsic routing mechanism on the Bench2Drive benchmark, examining the impact of the uncertainty threshold on driving performance, as well as the router’s accuracy and expert utilization.
@@ II-A End-to-End Autonomous Driving
End-to-end autonomous driving is an emerging paradigm that aims to directly map raw sensor inputs to vehicle control actions, thereby simplifying the system architecture and mitigating error propagation. Driving policies are predominantly learned via Imitation Learning (IL) to mimic expert demonstrations [5, 6, 7, 8, 9, 10, 11, 12] or Reinforcement Learning (RL) for dynamic policy optimization [16, 17, 18].
In 2022, Trajectory-guided Control Prediction (TCP) [11] was proposed as a simple and robust monocular vision baseline. This method innovatively integrates trajectory planning and direct control into a unified pipeline for joint learning and predictive fusion. Furthermore, techniques such as TransFuser [5] adopt the Transformer to fuse information from cameras and LiDAR. Beyond fusion, recent innovations also include vectorized scene representations, exemplified by VAD [7], which models driving scenes as fully vectorized elements to improve planning efficiency and robustness. Diffusion models have also emerged as a powerful tool, with DiffusionDrive [19] utilizing truncated diffusion policies to model multi-modal action distributions while simultaneously achieving real-time control.
@@ II-B Mixture-of-Experts in Autonomous Driving
The MoE architecture has emerged as a significant method for scaling LLMs and enhancing task specialization. In LLMs, sparse MoE designs boost model capacity and processing efficiency through conditional computation [15]. The strength of MoE lies in its ability to leverage individual experts’ strengths across varying data subsets or tasks, thereby improving overall model performance.
Despite promising results in LLMs, MoE’s application in end-to-end autonomous driving remains underexplored. Some existing studies have explored MoE architectures in autonomous driving for tasks such as rare scenario perception [20], long-tailed trajectory prediction[21], domain adaptation in different weathers [22], safe trajectory prediction and planning [23], continual adaptation to failure cases[24] and facilitating generalization of planner [25]. However, these existing approaches have not focused on leveraging MoE to enhance adaptive and robust performance in diverse scenarios.
Figure 3: A Conceptual Illustration of Representing Multimodal Policy Distribution with Mixture-of-Experts.
@@ III METHODOLOGY
Fig. 1 illustrates the overall architecture of our proposed GEMINUS framework. Our approach begins with the construction of a foundational end-to-end model, inspired by the design philosophies of TCP [11]. This model functions as a conventional single-expert baseline upon which we integrate our novel Dual-aware MoE architecture to form the complete GEMINUS framework.
@@ III-A Preliminaries
End-to-End Autonomous Driving. The objective of end-to-end autonomous driving is to directly map raw sensor inputs to corresponding trajectories or control actions. In this paper, the raw sensor input $x$ encompasses: a front-facing camera image $i$, the ego-vehicle speed $v$, a high-level navigation command $c$, and a goal point $(x_{g},y_{g})$. The raw sensor inputs are processed by the end-to-end model. Encoders first process these inputs to generate intermediate features. These features are then fed into a trajectory planner. The trajectory planner generates a planned trajectory that comprises waypoints in $K$ steps. This planned trajectory is then fed into a Proportional-Integral-Derivative (PID) controller. The controller subsequently produces the final longitudinal control signals: throttle $\in[0,1]$, brake $\in[0,1]$, and the lateral control signal: steer $\in[-1,1]$.
Mixture-of-Experts. The MoE architecture offers a principled approach to address the complexities of multimodal data distributions employing a ”divide and conquer” strategy [15]. Introducing the MoE Framework to end-to-end trajectory planning, the overall policy distribution $p_{\theta}(Y\mid X)$ is typically represented as a probabilistic mixture of policy distributions of components $K$, each parameterized by an expert $m_{\theta}(Y\mid Z=k,X)$ and weighted by a gating network $q_{\theta}(Z=k\mid X)$, formalized as:
@@
As conceptually illustrated in Fig. 3, this probabilistic formulation allows a model to tackle the mode averaging problem by representing the multi-modal distribution of driving policies as a mixture of multiple expert policy distributions[26, 27]. A common and effective simplification in deep learning implementations is to use a Hard Assignment approach, where a router selects a single, most suitable expert for a given input. This hard assignment strategy is highly effective as it avoids averaging the outputs of multiple experts, thus further mitigating mode averaging. Building upon these theoretical underpinnings, this paper proposes GEMINUS, a distinctive MoE framework that is specifically tailored for diverse and complex autonomous driving scenarios. At its core, a Dual-aware Router possesses scenario and uncertainty awareness to dynamically activate experts from a Global Expert and a Scene-Adaptive Experts Group. During inference, the Dual-aware Router processes intermediate features $x$ extracted by the encoders. It determines the final output $y$ based on the uncertainty measure $U(x)$ and the scores of scene experts $S_{E}(x)$, formalized as:
@@
Here $x$ is the input feature. $\tau$ denotes a predefined uncertainty threshold. $\mathcal{S}$ is the set of all the scene-adaptive experts. When $U(x)$ is low ($U(x)<\tau$), the expert with the highest routing score $S_{E_{i}}(x)$ is selected. This achieves precise and scenario-specific planning in feature-distinct scenarios. In contrast, when the uncertainty of the scenario $U(x)$ is high ($U(x)\geq\tau$), the model selects the Global Expert $f_{\text{global}}(x)$. This ensures robust performance in feature-ambiguous scenarios. This design allows GEMINUS to effectively avoid the mode averaging problem, thus achieving adaptive and robust performance in diverse scenarios.
@@ III-B Single-Expert Baseline
Feature Encoders. The image encoder is a ResNet34 architecture pretrained on ImageNet [28], which processes the view of the front-facing camera. Concurrently, a measurement encoder processes vehicle state information, including its speed $v$, the navigation command $c$, and the goal point $(x_{g},y_{g})$. The feature vectors from both encoders are then concatenated into a combined feature vector $F$ for the planner.
Trajectory Planner. The trajectory planner first processes the combined feature $F$ through linear layers for downsampling. The resulting vector is then fed into a GRU-based model [29] that autoregressively generates a sequence of waypoints for the next $K=4$ steps. Finally, these waypoints are used by downstream controllers to produce throttle, brake, and steer commands.
@@ III-C Scenario-aware Routing Mechanism
Vanilla MoE aims to balance expert usage across GPUs to utilize maximum benefit from input features. However, this leads to inefficient knowledge sharing among experts when dealing with heterogeneous input distributions. For example, driving policies for a Merging scenario differ significantly from an Emergency Brake scenario. To address this inefficiency and foster specialized knowledge, a scenario-aware routing mechanism is introduced. This mechanism draws inspiration from dataset-aware routing in[30].
@@
$\mathcal{L}_{\text{scenario}}=-\sum_{i=1}^{|E|}\mathbf{1}(h(s_{m})=i)\cdot\log p_{i}(x)$ (3)
$\mathbf{1}(\cdot)$ is the indicator function. This loss ensures that all inputs originating from a specific scenario category are primarily dispatched to their designated scene-adaptive expert. By selectively routing inputs based on their scenario feature, this mechanism promotes efficient knowledge specialization within each expert. This enables the model to learn adaptive driving policies.